# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 4.2 Communication: Pebble
In the following part we give a detailed description of our algorithm PerpExplore-Coloc-Pbl. Initially all the agents are in state Initial. In this state, an agent first declares its $Current-Node$ as $home$, after which initializes the variable $T_{time}=0$ (where, $T_{time}$ is the number of rounds passed since the agent has moved from state Initial) and gathers the ID of the remaining agents currently at $home$. Next, the agent with lowest ID, i.e., $a_{1}$ moves to state Leader, the agent with second lowest ID, i.e., $a_{2}$ moves to state Follower-Find, whereas the remaining agent, i.e., $a_{3}$ moves to state Backup. We now define an iteration for this algorithm. An iteration is defined to be a collection of $14n+$ 1 consecutive rounds starting from the latest round where all 3 agents along with 3 pebbles are at $home$ in state Initial. Note that, a new iteration fails to execute, if either at least one pebble or an agent gets destroyed by the black hole in the current iteration. More precisely, the meaning of failing an iteration implies that, an agent after ending its current iteration does not again start a new iteration by moving in to state Initial.
@@
Suppose, all the agents along with the pebbles successfully execute the $i$-th iteration and reaches $home$, i.e., neither any pebble nor any agent gets destroyed by the black hole. So, according to our algorithm, $a_{1}$ being the lowest ID, follows the following sequence $SEQ1$, whereas $a_{2}$ being the second lowest ID follows $SEQ2$, and the $a_{3}$ follows the sequence $SEQ3$. The above sequences are as follows:
@@
Execution of sequence $SEQ1$: $a_{1}$ performs this sequence, in which it first changes to state Leader, after finding it as the lowest ID agent at $home$ in state Initial. In state Leader, the agent performs the following checks: if the current node is $home$, then it checks that if the current node has the agent with second lowest ID, whereas the number of pebble at $home$ is 3 and $T_{time}\geq 2$. If all these conditions are satisfied, then the agent waits till $T_{time}=4n$ after which it moves to state Initial during $T_{time}=4n+1$. Otherwise, if it finds that it is with the second lowest ID agent and number of pebble at $home$ is also 3, but $T_{time}<2$, in this case the agent initializes $W_{time}=0$ (where $W_{time}$ is the waiting time of the agent) and then moves one hop along with the pebble to the next clockwise node. Otherwise, if the agent finds that neither it is with the second lowest ID agent nor the number of pebbles at $home$ is 3, then it waits for 5 rounds (i.e., waits when $W_{time}<6$), after which if still the above condition persists, then concludes $S_{lft}$ as the node which is at a clockwise distance of $n-2$ from $home$ and $S_{rgt}$ the node at a clockwise distance of $n-1$. Further, it updates $W_{time}=0$ and moves to the next state Detection.
@@
On the contrary, if the current node of $a_{1}$ is not $home$, then also it performs the following checks: first, it checks whether it is with the second lowest ID agent (i.e., $a_{2}$ in this case), if so, then initializes $W_{time}=0$ and then moves one hop clockwise to the next node along with the pebble it is accompanying. Otherwise, if it is not with the second lowest ID agent, then waits for 3 rounds (i.e., waits when $W_{time}<4$), after which if still it is without the second lowest ID agent, then the agent moves to state Report-Leader, while leaving the pebble at the current node and moving one hop to the clockwise node.
@@
Note that until and unless there is no anomalies detected by the agent $a_{1}$, it ends an iteration, by changing to state Initial from Leader. Otherwise, if any anomaly is detected, then $a_{1}$ ends the current iteration by either changing to state Detection or Report-Leader from the state Leader, in which case, it never changes to state either Leader or Initial.
@@
Execution of sequence $SEQ2$: $a_{2}$ being the second lowest Id agent in state Initial, performs this sequence in an iteration. On successful completion of this iteration, an agent starting from Initial, performs the sub-sequence ${Follower-SEQ}$, $n$ times and then again changes to state Initial. The sub-sequence ${Follower-SEQ}$ symbolizes the sequence Follower-Find $\rightarrow$ Follower-Collect $\rightarrow$ Follower-Find, which an agent performs $n$ times until it again changes to state Initial (i.e., ${Follower-SEQ}$ is denoted by ${Follower-SEQ}^{n}$). While executing ${Follower-SEQ}$, the agent in state Follower-Find checks, whether the current node does not contain the lowest ID agent and the $Move$ parameter is also set to 0 (the parameter $Move$ is either 0 or 1, when 1 it symbolizes that the current node must contain the lowest ID agent, when 0, it means the current node is the adjacent counter clockwise node with respect to the node which contains the lowest ID agent), if so then it moves one hop clockwise leaving the pebble at the current node, while updating $Move$ to 1. Otherwise, if it finds $Move$ to be 1 whereas the current node does not contain the lowest ID agent, then it stops the current iteration, detects that current node is black hole and starts performing perpetual exploration, avoiding the black hole node. On the other hand, if the current node contains the lowest ID agent, then $a_{2}$ directly changes to state Follower-Collect. In state Follower-Collect, $a_{2}$ first moves one hop counter-clockwise, then in the current node either finds a pebble or not. If it finds a pebble, then moves one hop clockwise along with the pebble. Now, if $T_{time}=4n+1$ and current node is $home$, then stops performing $Follower-SEQ$ and moves to state Initial. Otherwise, updates $Move$ to 0 and changes to state Follower-Find. If the pebble is not found, then it concludes current node to be the black hole and starts performing perpetual exploration avoiding this node.
@@
Note that, the agent only changes to state Initial only when it reaches $home$ again, which is after it performs ${Follower-SEQ}$ for $n$ times starting from $SEQ2$. If at any point, if $a_{2}$ finds any irregularities, based on whether a pebble is present or not, then only it directly concludes the black hole position. In all other cases, it follows the ${Follower-SEQ}$ sequence and changes to state Follower-Collect.
@@
  • The current node, i.e., $home$ has only the lowest and highest ID agents, i.e., $a_{1}$ and $a_{3}$, whereas the number of pebble at $home$ is 1. If this condition is satisfied, then the agent changes its state to Find-Pebble. Note that, this condition is satisfied only when $a_{2}$ fails to reach $home$ while performing this iteration, i.e., enters the black hole, while $a_{1}$ has detected some anomalies, which instigates it to leave the pebble at the current position at which it detects the anomaly, and henceforth moves back to $home$ while in state Report-Leader.
@@
  • $home$ has only the highest and lowest ID agents, i.e., $a_{1}$ and $a_{3}$, but number of pebble at $home$ is 2, which implies that $a_{2}$ has failed to return to $home$ even when $T_{time}=4n$. In this case, $a_{3}$ concludes $S_{lft}$ to be the node which is at a clockwise distance of $n-2$ from $home$, whereas $S_{rgt}$ to be the node which is at a clockwise distance of $n-1$ from $home$. Finally, it updates $W_{time}=0$ and then moves to state Detection.
@@
  • $home$ has all the three agents and moreover, number of pebbles present at $home$ is also 3. In this case, $a_{3}$ changes to state Initial, completing the iteration without detecting any anomalies.
@@
  • After $4n$ rounds, there is no other agent except $a_{3}$ present at $home$. In this case, the agent move one hop, clockwise without the pebble present at $home$, and then changes to state Find-BH. This case arises only when both the agents $a_{1}$ and $a_{2}$ fail to reach $home$, while performing this iteration.
@@
Now, next we define the states Report-Leader, Find-Pebble, Find-BH, Detection. An agent moves to either of these states only when it detects some anomalies.
@@
The state Report-Leader, is executed by only the lowest ID agent, i.e., $a_{1}$, while it is in state Leader. The anomaly detected is as follows, $a_{1}$, after reaching a new node that is not $home$, along with the pebble it is carrying, waits for at most 3 rounds, for second lowest ID agent, i.e., $a_{2}$ to arrive (which is in state Follower-Find). If $a_{2}$ does not arrive even after the waiting, $a_{1}$ concludes that the agent (i.e., $a_{2}$) has entered the black hole, this triggers $a_{1}$ to move to state Report-Leader. In this state, it moves clockwise until it finds the highest ID agent, i.e., $a_{3}$ at $home$. Then waits until $T_{time}=4n$, after which changes state to Find-Pebble.
@@ 3 In State Initial:
18       if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}\geq 2$  then
@@
22       else if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ Id\wedge\#Pebble=3\wedge T_{time}<2$ then
@@ 31                   $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
54       Detect the $Current-Node$ as black hole and continue exploring the ring perpetually avoiding this node.
@@
69       Detect the $Current-Node$ as black hole and continue perpetually exploring the ring avoiding this node.
@@ 54 In State Report-Leader:
63 if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=1$ then
@@
66 else if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=2$ then
@@ 68       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
71 else if $Current-Node$ has both lowest and second lowest Id agent and #$\#Pebble=3$ then
@@
75       Move in a clockwise direction without pebble and change to state Find-BH.
@@
81       Conclude the next node along clockwise direction is the black hole, and continue perpetual exploration avoiding the black hole node.
@@
88 Move counter-clockwise without the pebble.
@@ 92       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
102       If other agent is there it repeats the same procedure as stated in line 86 of Algorithm 2. Otherwise detects $S_{rgt}$ as the black hole and start exploring avoiding that node.
@@
105       If other agent is there it repeats the same procedure as stated in line 90 of Algorithm 2. Otherwise detects $S_{lft}$ as the black hole and start exploring avoiding that node.
The state Find-Pebble, is performed together by only $a_{1}$ and $a_{3}$, i.e., the lowest and highest ID agent. This state can be reached by $a_{1}$ only via Report-Leader state, whereas by $a_{3}$ from the state Backup. The main idea for the agents in this state is to move counter-clockwise till they find the pebble left by $a_{1}$ when it detects some anomaly (i.e., $a_{2}$ has entered the byzantine black hole) and moved into state Report-Leader. This pebble acts as a marker which indicates that black hole is either the counter-clockwise neighbour of this node or, it is at a counter clockwise distance of 2 hop from it. This node at which it finds the pebble is declared to be the new $home$, from which they conclude $S_{lft}$ and $S_{rgt}$ to be the nodes which are at a clockwise distance of $n-2$ and $n-1$, respectively from this new $home$. Finally, they updates $W_{time}=0$ and then changes to state Detection. Note that, this state is performed by an agent during the last round of the current iteration, (i.e., when $T_{times}=4n+1$).
The state Find-BH, is only executed by the highest ID agent, i.e., $a_{3}$. This state is executed by $a_{3}$, only when it finds no other agent at $home$, at the last round of the current iteration. This situation can only occur, when both the agents $a_{1}$ and $a_{2}$ has entered the black hole. So, this triggers $a_{3}$ to change to state Find-BH, in which it moves along a clockwise direction, until a pebble is found. Whenever, a pebble is found, it concludes the next node to be the black hole, and starts performing perpetual exploration avoiding this node.
Finally, the state Detection, is performed by an agent (i.e., $a_{1}$ or $a_{3}$) only if they know the position of $S_{lft}$ and $S_{rgt}$, which are not only two consecutive nodes in a suspicious region $S$. Moreover, these two nodes are exactly at a distance of clockwise $n-1$ and $n-2$ distance from $home$ (the $home$ can be the initial starting node, or it can be the updated $home$ as well). In this state, the agent with lowest ID, i.e., $a_{1}$, moves clockwise till it reaches $S_{rgt}$, and then again returns back to $home$. After returning back, if it finds $a_{3}$, then again it performs the same procedure. On the contrary, if it does not find $a_{3}$, then it understands that $S_{rgt}$ is the black hole node, and hence continues perpetual exploration avoiding this node.
On the other hand, $a_{3}$ in this state moves counter-clockwise direction from $home$, i.e., reaches $S_{lft}$, after which it returns back to $home$ again following the same path and then waits for $62n-$ 6 rounds at $home$ and then checks for the other agent, i.e., $a_{1}$. If it finds $a_{1}$, then again continues to perform the same movement, otherwise it concludes that $S_{lft}$ is the black hole node, and hence starts the perpetual exploration of the ring just by avoiding this black hole node.
4.2.1 Correctness and Complexity
In this section we discuss the correctness and complexity of our algorithm PerpExplore-Coloc-Pbl.

### v2

@@ 9.1 Detailed Description and Pseudocode
In the following part, we give a detailed description of our algorithm PerpExplore-Coloc-Pbl. Initially, all the agents are in state Initial. In this state, an agent first declares its $Current-Node$ as $home$, after which initializes the variable $T_{time}=0$ (where, $T_{time}$ is the number of rounds passed since the agent has moved from state Initial) and gathers the ID of the remaining agents currently at $home$. Next, the agent with the lowest ID, i.e., $a_{1}$ moves to state Leader, the agent with the second lowest ID, i.e., $a_{2}$ moves to state Follower-Find, whereas the remaining agent, i.e., $a_{3}$ moves to state Backup. We now define an iteration for this algorithm. An iteration is defined to be a collection of $14n+$ 1 consecutive rounds starting from the latest round where all 3 agents along with 2 pebbles are at $home$ in state Initial. Note that, a new iteration fails to execute if either at least one pebble or an agent gets destroyed by the Byzantine black hole in the current iteration. More precisely, the meaning of failing an iteration implies that an agent after ending its current iteration does not again start a new iteration by moving into state Initial.
@@
Suppose, all the 3 agents along with 2 pebbles successfully execute the $i$-th iteration and reach $home$, i.e., neither any pebble nor any agent gets destroyed by the Byzantine black hole. In this situation, according to our algorithm, $a_{1}$ being the lowest ID, follows the following sequence $SEQ1$, whereas $a_{2}$ being the second lowest ID follows $SEQ2$, and the $a_{3}$ follows the sequence $SEQ3$. The above sequences are as follows:
@@
Execution of sequence $SEQ1$: $a_{1}$ performs this sequence, in which it first changes to state Leader, after finding it as the lowest ID agent at $home$ in state Initial. In state Leader, the agent performs the following checks: if the current node is $home$, then it checks that if the current node has the agent with the second lowest ID, whereas the number of pebbles at $home$ is 2 and $T_{time}\geq 2$. If all these conditions are satisfied, then the agent waits till $T_{time}=4n$ after which it moves to state Initial during $T_{time}=4n+1$. Otherwise, if it finds that it is with the second lowest ID agent and the number of pebbles at $home$ is also 2, but $T_{time}<2$, in this case, the agent initializes $W_{time}=0$ (where $W_{time}$ is the waiting time of the agent) and then moves one hop along with the pebble to the next clockwise node. Otherwise, if the agent finds that neither it is with the second lowest ID agent nor the number of pebbles at $home$ is 2, then it waits for 5 rounds (i.e., waits when $W_{time}<6$), after which if still the above condition persists, then concludes $S_{lft}$ as the node which is at a clockwise distance of $n-2$ from $home$ and $S_{rgt}$ the node at a clockwise distance of $n-1$. Further, it updates $W_{time}=0$ and moves to the next state Detection.
@@
On the contrary, if the current node of $a_{1}$ is not $home$, then also it performs the following checks: first, it checks whether it is with the second lowest ID agent (i.e., $a_{2}$ in this case), and if so, then initializes $W_{time}=0$ and then moves one hop clockwise to the next node along with the pebble it is accompanying. Otherwise, if it is not with the second lowest ID agent, then wait for 3 rounds (i.e., waits when $W_{time}<4$), after which if still it is without the second lowest ID agent, then the agent moves to state Report-Leader while leaving the pebble (it is accompanying) at the current node and moving one hop to the clockwise node.
@@
Note that until and unless there are no anomalies detected by the agent $a_{1}$, it ends an iteration, by changing to state Initial from Leader. Otherwise, if any anomaly is detected, then $a_{1}$ ends the current iteration by either changing to state Detection or Report-Leader from the state Leader, in which case, it never changes to state either Leader or Initial.
@@
Execution of sequence $SEQ2$: $a_{2}$ being the second lowest ID agent in state Initial, executes this sequence in an iteration. On successful completion of this iteration, an agent starting from Initial, performs the sub-sequence ${Follower-SEQ}$, $n$ times and then again changes to state Initial. The sub-sequence ${Follower-SEQ}$ symbolizes the sequence Follower-Find $\rightarrow$ Follower-Collect $\rightarrow$ Follower-Find, which an agent performs $n$ times until it again changes to state Initial (i.e., ${Follower-SEQ}$ is denoted by ${Follower-SEQ}^{n}$). While executing ${Follower-SEQ}$, the agent in state Follower-Find checks, whether the current node does not contain the lowest ID agent, and the $Move$ parameter is also set to 0 (the parameter $Move$ is either 0 or 1, when 1 it symbolizes that the current node must contain the lowest ID agent, when 0, it means the current node is the adjacent counter-clockwise node with respect to the node which contains the lowest ID agent), if so then it moves one hop clockwise leaving the pebble at the current node, while updating $Move$ to 1. Otherwise, if it finds $Move$ to be 1 whereas the current node does not contain the lowest ID agent, then it stops the current iteration, detects that the current node is a Byzantine black hole, and starts performing perpetual exploration, avoiding the Byzantine black hole node. On the other hand, if the current node contains the lowest ID agent, then $a_{2}$ directly changes to state Follower-Collect. In state Follower-Collect, $a_{2}$ first moves one hop counter-clockwise, then in the current node either finds a pebble or not. If it finds a pebble, then moves one hop clockwise along with the pebble. Now, if $T_{time}=4n+1$ and the current node is $home$, then stops performing $Follower-SEQ$ and moves to state Initial. Otherwise, updates $Move$ to 0 and changes to state Follower-Find. If the pebble is not found, then it concludes the current node to be the Byzantine black hole and starts performing perpetual exploration avoiding this node.
@@
Note that, the agent only changes to state Initial only when it reaches $home$ again, which is after it performs ${Follower-SEQ}$ for $n$ times starting from $SEQ2$. If at any point, $a_{2}$ finds any irregularities, based on whether a pebble is present or not, then only it directly concludes the Byzantine black hole position. In all other cases, it follows the ${Follower-SEQ}$ sequence and changes to state Follower-Collect.
@@
  • The current node, i.e., $home$ has only the lowest and highest ID agents, i.e., $a_{1}$ and $a_{3}$, whereas the number of pebble at $home$ is 0. If this condition is satisfied, then the agent changes its state to Find-Pebble. Note that, this condition is satisfied only when $a_{2}$ fails to reach $home$ while performing this iteration, i.e., enters the Byzantine black hole, while $a_{1}$ has detected some anomalies, which instigates it to leave the pebble at the current position at which it detects the anomaly, and henceforth moves back to $home$ while in the state Report-Leader.
@@
  • $home$ has only the highest and lowest ID agents, i.e., $a_{1}$ and $a_{3}$, but the number of pebbles at $home$ is 1, which implies that $a_{2}$ has failed to return to $home$ even when $T_{time}=4n$. In this case, $a_{3}$ concludes $S_{lft}$ to be the node which is at a clockwise distance of $n-2$ from $home$, whereas $S_{rgt}$ to be the node which is at a clockwise distance of $n-1$ from $home$. Finally, it updates $W_{time}=0$ and then moves to state Detection.
@@
  • $home$ has all three agents and the number of pebbles present at $home$ is 2. In this case, $a_{3}$ changes to state Initial, completing the iteration without detecting any anomalies.
@@
  • After $4n$ rounds, there is no other agent except $a_{3}$ present at $home$. In this case, the agent moves one hop clockwise, and then changes to state Find-BH. This case arises only when both the agents $a_{1}$ and $a_{2}$ fail to reach $home$ while performing this iteration.
@@
Next we define the states Report-Leader, Find-Pebble, Find-BH, Detection. An agent moves to either of these states only when it detects some anomalies.
@@
The state Report-Leader, is executed by only the lowest ID agent, i.e., $a_{1}$, while it is in state Leader. The anomaly detected is as follows, $a_{1}$, after reaching a new node that is not $home$, along with the pebble it is carrying, waits for at most 3 rounds, for the second lowest ID agent, i.e., $a_{2}$ to arrive (which is in state Follower-Find). If $a_{2}$ does not arrive even after the waiting, $a_{1}$ concludes that the agent (i.e., $a_{2}$) has entered the Byzantine black hole, this triggers $a_{1}$ to move to state Report-Leader. In this state, it moves clockwise until it finds the highest ID agent, i.e., $a_{3}$ at $home$. Then waits until $T_{time}=4n$, after which it changes its state to Find-Pebble.
The state Find-Pebble, is performed together by only $a_{1}$ and $a_{3}$, i.e., the lowest and highest ID agent. This state can be reached by $a_{1}$ only via Report-Leader state, whereas by $a_{3}$ from the state Backup. The main idea for the agents in this state is to move counter-clockwise till they find the pebble left by $a_{1}$ when it detects some anomaly (i.e., $a_{2}$ has entered the Byzantine black hole) and move into state Report-Leader. This pebble acts as a marker which indicates that the Byzantine black hole is either the counter-clockwise neighbor of this node or, it is at a counter-clockwise distance of 2 hops from it. This node at which it finds the pebble is declared to be the new $home$, from which they conclude $S_{lft}$ and $S_{rgt}$ to be the nodes which are at a clockwise distance of $n-2$ and $n-1$, respectively from this new $home$. Finally, they update $W_{time}=0$ and then change to state Detection. Note that, this state is performed by an agent during the last round of the current iteration, (i.e., when $T_{times}=4n+1$).
The state Find-BH, is only executed by the highest ID agent, i.e., $a_{3}$. This state is executed by $a_{3}$, only when it finds no other agent at $home$, at the last round of the current iteration. This situation can only occur, when both the agents $a_{1}$ and $a_{2}$ have entered the Byzantine black hole. So, this triggers $a_{3}$ to change to state Find-BH, in which it moves along a clockwise direction until a pebble is found. Whenever a pebble is found, it concludes the next node to be the Byzantine black hole, and starts performing perpetual exploration avoiding this node.
Finally, the state Detection, is performed by an agent (i.e., $a_{1}$ or $a_{3}$) only if they know the position of $S_{lft}$ and $S_{rgt}$, which are not only two consecutive nodes in a suspicious region $S$. Moreover, these two nodes are exactly at a distance of clockwise $n-1$ and $n-2$ distance from $home$ (the $home$ can be the initial starting node, or it can be the updated $home$ as well). In this state, the agent with the lowest ID, i.e., $a_{1}$, moves clockwise till it reaches $S_{rgt}$, and then again returns back to $home$. After returning back, if it finds $a_{3}$, then again it performs the same procedure.
@@ 3 In State Initial:
18       if $with\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ ID\wedge\#Pebble=2\wedge T_{time}\geq 2$  then
@@
22       else if $With\leavevmode\nobreak\ Second\leavevmode\nobreak\ Lowest\leavevmode\nobreak\ ID\wedge\#Pebble=2\wedge T_{time}<2$ then
@@ 31                   $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
54       Detect the $Current-Node$ as Byzantine black hole and continue exploring the ring perpetually avoiding this node.
@@
69       Detect the $Current-Node$ as Byzantine black hole and continue perpetually exploring the ring avoiding this node.
@@ 54 In State Report-Leader:
63 if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=0$ then
@@
66 else if $Current-Node$ has only the lowest and highest ID agents and #$\#Pebble=1$ then
@@ 68       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
71 else if $Current-Node$ has both lowest and second lowest ID agent and #$\#Pebble=2$ then
@@
75       Move in a clockwise direction and change to state Find-BH.
@@
81       Conclude the next node along clockwise direction is the Byzantine black hole, and continue perpetual exploration avoiding the Byzantine black hole node.
@@
88 Move in a counter-clockwise direction.
@@ 92       $S_{rgt}=$ node at a clockwise distance of $n-1$ from $home$
102       If other agent is there it repeats the same procedure as stated in line 86 of Algorithm 1. Otherwise detects $S_{rgt}$ as the Byzantine black hole and start exploring avoiding that node.
@@
105       If other agent is there it repeats the same procedure as stated in line 90 of Algorithm 1. Otherwise detects $S_{lft}$ as the Byzantine black hole and start exploring avoiding that node.
On the contrary, if it does not find $a_{3}$, then it understands that $S_{rgt}$ is the Byzantine black hole node, and hence continues perpetual exploration avoiding this node.
On the other hand, $a_{3}$ in this state moves counter-clockwise direction from $home$, i.e., reaches $S_{lft}$, after which it returns back to $home$ again following the same path and then waits for $62n-$ 6 rounds at $home$ and then checks for the other agent, i.e., $a_{1}$.
If it finds $a_{1}$, then again continues to perform the same movement, otherwise, it concludes that $S_{lft}$ is the Byzantine black hole node, and hence starts the perpetual exploration of the ring just by avoiding this Byzantine black hole node.
9.2 Correctness and Complexity
In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Coloc-Pbl.